OpenStudy (anonymous):

let f(x) be an even function and g(x0 be an odd function. use the definition of even and odd functions to determine whether the functions are even odd or neither . 1. F(x)*G(x) 2. F(x)/G(X) 3. F(G(x)) 4. G(F(x)) 5. F(x)+G(x) 6. F(x)-G(x)

OpenStudy (unklerhaukus):

f(x) is an even function ⟹ f(-x) = f(x) g(x) is an odd function ⟹ g(-x) = -g(x)

OpenStudy (unklerhaukus):

1. f(x)*g(x) to find if this product of functions is odd or even compute f(-x)*g(-x) and see if is equal or opposite to f(x)*g(x)

OpenStudy (unklerhaukus):

@Reggie97g

OpenStudy (anonymous):

i'm very confused on how you know if its odd or even . is it like positive or negative ?

OpenStudy (unklerhaukus):

Some function A(x) is even, means that A(-x) = A(x) Some function B(x) is odd, means that B(-x) = -B(x)

OpenStudy (unklerhaukus):

It might be more clear if , for the first question 1. f(x)*g(x) we set h(x) = f(x)*g(x) and then compute h(-x), in-terms of f(x) and g(x) if we find that h(-x) = h(x) then we have shown h(x) is even if we find that h(-x) = -h(x) then we have shown h(x) is odd

OpenStudy (unklerhaukus):

h(x) = f(x)*g(x) h(-x) = f(-x)*g(-x) = ( use the fact that f(x) is even, and g(x) is odd ) =

OpenStudy (unklerhaukus):

can you work out the next line @Reggie97g ?

OpenStudy (anonymous):

h(x) = f(x)*-g(x) ?

OpenStudy (unklerhaukus):

h(-x) = f(-x)*g(-x) you mean h(-x) = f(x) * -g(x) right?

OpenStudy (anonymous):

oops yes

OpenStudy (unklerhaukus):

so is h(-x) equal or opposite to h(x) ?

OpenStudy (anonymous):

it is opposite right ?

OpenStudy (anonymous):

no no its equal i just saw what you put sorry

OpenStudy (unklerhaukus):

h(-x) = f(x) * -g(x) = -f(x)*g(x) = -h(x) h(-x) is opposite to h(x)

OpenStudy (anonymous):

but before you put h(-x)=h(x) !

OpenStudy (anonymous):

ooh that shows its even and not odd

OpenStudy (unklerhaukus):

is h(x) even or odd?

OpenStudy (anonymous):

even

OpenStudy (unklerhaukus):

how did you get that?

OpenStudy (anonymous):

if h(-x) = -h(x) then its odd and thats what the problem equals . so the function must be odd

OpenStudy (unklerhaukus):

thats better

OpenStudy (unklerhaukus):

For the next one h(x) = f(x) / g(x) h(-x) = f(-x) / g(-x) = (simplify this )

OpenStudy (unklerhaukus):

(remember, you know f(x) is even, and you also know g(x) is odd)

OpenStudy (anonymous):

= f(x) / -g(x) = -h(x) so its odd

OpenStudy (unklerhaukus):

great!, so far you have proved that 1. the product of an even function and an odd function is and odd function 2. the quotient of an even function and an odd function is and odd function

OpenStudy (unklerhaukus):

3. is the composition of functions this time h(x) = f(g(x)) = f ∘ g (x) the method is the same find h(-x)

OpenStudy (unklerhaukus):

*an